Find the future value on $5,000 in three years if the interest is compounded semiannually at 12%.
Answer by TMils(8)

Find the future value on $5,000 in three years if the interest is compounded semiannually at 12%.
semiannually = 2x/yr
A=P(1+r/n)^nt
A = 5000(1+(.12/2))^2x3
A = 5000(1.06)^6
A= $7092.60 (7092.595561 rounded up to the nearest 2 decimal places)